I have tried to figure this out so that I can help my daughter but I just can not remember how to do it. She didn't understand it so maybe you can help me help her. ok, here's the question, A rectangular prism has a volume of 54 cubic feet. If the length is doubled, the width is tripled, and the height is quadrupled, what is the new volume of the prism. Hope you can help. Thank You, Jennifer
Answer by ankor@dixie-net.com(22740)

A rectangular prism has a volume of 54 cubic feet. If the length is doubled, the width is tripled, and the height is quadrupled, what is the new volume of the prism.
:
It would be (2 * 3 * 4) * 54 = 1296 cu ft
:
You can prove this: Assume the original is 6 * 3 * 3 (which is 54)
The new size:
2(6) * 3(3) * 4(3) =
12 * 9 * 12 = 1296